Circular No. 8370

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004da IN NGC 6901 T. Boles, Coddenham, England, reports the discovery of an appar-

ent supernova (mag 15.2) on unfiltered CCD images taken on July 6.099 and 7.031 UT with a 0.35-m reflector. The new object is located at α = 20^{h}22^{m}21^{s}.24, δ = +6^{o}25′50′′.3, which is ≈ 5′′.4 west and 2′′.3 north of the center of NGC 6901. SN 2004da is not present on Boles’ images from May 5 (limiting mag 19.5) and June 14, and it is not present on Digitized Sky Survey red (1990 Aug. 19) or blue (1988 July 8) plates.

V5113 SAGITTARII AND V2574 OPHIUCHI Further to IAUC 8368, R. J. Rudy, D. K. Lynch, S. Mazuk, C. C.

Venturini, R. C. Puetter, and R. B. Perry report 0.4–2.5-μm NIRIS spectroscopy of V5113 Sgr on June 24.06 UT, ≈ 9 months after outburst (cf. IAUC 8204, 8206). This Fe II-type nova has reached moderately high excitation, displaying strong He II lines as well as the coronal lines [S VIII] 0.9911-μm and [Si VI] 1.9641-μm. The unidentified novae lines at 1.1114, 1.1900, 1.5545, and 2.0996 μm were also present. The nebular [O III] lines were very strong, as was the [S III] feature at 0.9532 μm. The continuum shows a sharp rise beyond 2 μm, indicative of hot dust.

Puetter, Rudy, Lynch, Mazuk, Venturini, and Perry add that their NIRIS observations of V2574 Oph (cf. IAUC 8323, 8324) on June 23.4 UT show that the nova is still in its low-excitation phase, showing lines of C I, N I, and Fe II. No He II or coronal lines are seen. The Balmer, Paschen, and Bracket lines are strong, as are the Lyβ-fluoresced O I lines at 0.8446 and 1.1287 μm. The lines are approximately gaussian in shape but are asymmetric near their peaks (FWHM ≈ 1500 km/s). The slow development of this nova suggests that continued monitoring should be possible for several years.

PSR 0833–45 R. Dodson, Institute of Space and Astronautical Science; S. Buchner,

Hartebeesthoek Radio Astronomy Observatory (HartRAO); B. Reid and D. Lewis, University of Tasmania; and C. Flanagan, Johannesburg Planetarium, report observing a sudden spin-up (“glitch”) in the Vela pulsar (cf. IAUC 7347) on July 7.09 UT. Preliminary model fitting gives a fractional decrease of 2.1 × 10^{−6} in the rotation period. The jump was first detected by the HartRAO 26-m telescope and observed at three frequencies (635.034, 990.25, and 1390.64 MHz) with the Mount Pleasant 14-m dish.

2004 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green